if m be the slope of a  tangent  to the curve ey=1+x2 then  find  mode of m. 

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

ey=1+x^2,          ey(dy/dx)=2x,(1+x2)(dy/dx)=2x,(dy/dx)=(2x/(1+x2), ((y-loge(1+xo^2))/(x-xo))=(2x/(1+x^2))

((loge(1+x^2)-loge(1+xo^2)))/(x-xo)= (2x/(1+x2)), solve for x. An algebraic(non-numerical), closed form solution might be IMPOSSIBLE because the equation in x is a TRANSCENTAL EQUATION.

Shounak Anand
6 years ago
|m|